Eight years ago, Cindy was twice as old as Sally. Right now, Cindy is five years older than Sally. Find their current ages.
Answer by ikleyn(52778)

.
Eight years ago, Cindy was twice as old as Sally. Right now, Cindy is five years older than Sally. Find their current ages.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

C = S + 5,         (1)    ("Cindy is five years older than Sally.")
C - 8 = 2*(S-8).   (2)    ("Eight years ago, . . . ")


Substitute (1) into (2). You will get


(S + 5) - 8 = 2*(S-8),

S - 3 = 2S - 16,

2S - S = 16 - 3,


S = 13.

Answer.  Sally is 13, Cindy is 13+5 = 18.

By the way, the phrase "Right now, Cindy is five years older than Sally." sounds strange for the human ear.

If "Cindy is five years older than Sally right now", she always IS five years older than Sally until both are alive.
------------------------------


There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".